OpenStudy (anonymous):

How the heck would I simplify this??? (1/cot^2(theta))+sec(theta)cos(theta)

OpenStudy (anonymous):

Using the identities cot = 1/tan(x) and sec(x) = 1/cos(x).

OpenStudy (mertsj):

Well you could start by recognizing that sec and cos are reciprocals and so their product is 1

OpenStudy (anonymous):

\[\frac{ 1 }{ \cot^2 \theta } \sec(\theta)\cos(\theta) \]

OpenStudy (anonymous):

I knew those did that, but would I have to multiply th fraction by.. tangent or something?

OpenStudy (solomonzelman):

These 3 identities might help.\[\frac{1}{Cos^2θ} =Sec^2θ,~~~~~~~~~~~~~~\frac{1}{Sec^2θ} =Cos^2θ,~~~~~~~~~~Cosθ~Secθ=1\]

OpenStudy (solomonzelman):

Oh, cot, my bad

OpenStudy (mertsj):

The tangent is the reciprocal of the cotangent so 1/cot^2=tan^2

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{1}{Cotθ} ~Secθ~Cosθ } \] \[\huge\color{blue}{ Tanθ~Secθ~Cosθ } \] \[USE~~~THIS~~~IDENTITY:~~~~~~~~~~~~~~~~~\huge\color{red}{ Secθ~Cosθ =1~~} \]

OpenStudy (anonymous):

I can't see that, but one of my choices is tan^2(theta). So that's right, right? None of them have +1.

OpenStudy (solomonzelman):

Yes, you are right.

OpenStudy (solomonzelman):

Also, in my work, I made an error, it's cot^2θ and tan^2θ

OpenStudy (mertsj):

So now you have reduced it to tan^2+1. You have an identity for that.

OpenStudy (anonymous):

Yay. You're fine. I saw it. :) I have a few more. Do you want me to post it seperately?